OpenStudy (anonymous):

What is the solution to the equation 14 x – 3 = 21 ?

OpenStudy (johnweldon1993):

So your ultimate focus...is to solve for 'x' 14x - 3 = 21 You need to isolate it....so add 3 to both sides of this equation 14x - 3 = 21 +3 +3 --------------- 14x = 24 Now to solve for 'x'...you divide both sides by 14... x = 24/14 x = ?

OpenStudy (the_fizicx99):

Add 3 to both sides, divide both sides by 14 and that x. Plug that in for x and subtract 3 14(?) - 3 = 21

OpenStudy (anonymous):